OpenStudy (anonymous):

In the problem(5a^2b)(-6a^3b^4) , Terrence got -30a^6b^4. Find his 2 mistakes and correct them

OpenStudy (anonymous):

To multiply powers add the exponents.

OpenStudy (anonymous):

\[5a^{2b}-\left( 6a^{3b} \right)^{4}\] This?

OpenStudy (anonymous):

oops, multiplied together, not subtracted

OpenStudy (anonymous):

\[\left( 5a^{2}b \right)\left( -6a^3b^4 \right)\] When exponenets with the same base are multiplied together, you add their exponents. So in this case, a to the 2nd is multiplied by a to the 3rd, so you're left with a to the 5th. b is multipied by b to the 4th, so you're left with b to the 5th. 5 and -6 multiply to -30 and stay in front. Your correct answer is \[-30a^5b^5\]